# Boosting Adverse Weather Crowd Counting via Multi-queue Contrastive Learning

### v2

@@ 5 Conclusion and Limitations
In this paper, we propose a contrastive learning-based method called MQCL to tackle the problem of class-imbalanced adverse weather crowd counting and synthesize a new adverse weather crowd counting dataset. The core contributions of this paper are multi-queue MoCo and the contrastive-learning-based refining method, which respectively achieve weather-aware representation learning and representation conversion, addressing the dual challenges of image degradation and class imbalance. Our work has achieved significant performance improvements compared to the baseline but there are still some limitations. For example, we only utilized synthetic methods to construct the dataset, leading to a certain domain gap between real-world weather. Additionally, for this study, our focus is on the application of multi-queue MoCo in the adverse weather crowd counting task. But it is a generic method to deal with data imbalance and is not only applicable to the crowd counting task but also holds significant reference value for other domains. Its performance in other tasks awaits future research.
6 Acknowledgement
This work was partially supported by the National Natural Science Foundation of China (62476130, 62176123), and Research Fund of Guangxi Key Lab of Multi-source Information Mining & Security, China (MIMS23-M-03).

### v3

@@ 5 Conclusion and Limitations
In this paper, we propose a contrastive learning-based method called MQCL to tackle the problem of class-imbalanced adverse weather crowd counting. The core contributions of this paper are multi-queue MoCo and the contrastive-learning-based refining method, which respectively achieve weather-aware representation learning and representation conversion, addressing the dual challenges of image degradation and class imbalance. Through the proposal of MQCL, we aim to provide a novel multi-domain perspective for addressing the crowd counting task. Nevertheless, our work still has limitations. For example, our focus is on the application of multi-queue MoCo in the crowd counting task. But it is a generic method to deal with data imbalance and holds significant reference value for other domains. Exploring its potential in more tasks is a direction for our future work.